# How to Correctly Report LLM-as-a-Judge Evaluations

### v2

@@ 9 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Adaptive Allocation of Calibration Samples
We propose an adaptive procedure for allocating calibration samples between the two ground-truth classes (‘correct’ and ‘incorrect’) using a small pilot calibration set. Leveraging pilot estimates of sensitivity $\tilde{q}_{1}$ and specificity $\tilde{q}_{0}$ in Algorithm 1, the algorithm attains an optimal allocation under a fixed total calibration budget, leading to short confidence intervals.
Algorithm 1 Adaptive allocation of calibration samples
@@ B.1 Deriving the Variance of Estimators
We now derive the asymptotic variance of $\hat{\theta}$ using the delta method (dorfman1938note; ver2012invented) for $\hat{\theta}=\frac{\hat{p}+\hat{q}_{0}-1}{\hat{q}_{0}+\hat{q}_{1}-1}$. The first order derivatives with respect to $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$ are
@@ B.2 Proofs of Propositions
Suppose that $m:=2m_{0}=2m_{1}$ and that $q:=q_{0}=q_{1}$ with $0.5<q\leq 1$. For sufficiently large $m\gtrsim 2q/(2q-1)^{2}$, the absolute bias of $\hat{\theta}$ in (2) is always smaller than that of $\hat{p}$ in (3) for all $\theta$.
@@
Hence, for sufficiently large $m$ satisfying $m\gtrsim 2q/(2q-1)^{2}$, we conclude the following for all $\theta$:
@@
Suppose that $\tilde{q}_{0}$ and $\tilde{q}_{1}$ are close to $1$, and let $\kappa:=(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$. Then the minimum length of the confidence interval defined in (9) is achieved when $\tilde{m}_{0}\approx(1/\tilde{p}-1)\sqrt{\kappa}\cdot\tilde{m}_{1}$.
Proof.
The length of the confidence interval in (9) is given by
$2z_{\alpha}\sqrt{\frac{\tilde{p}(1-\tilde{p})/\tilde{n}+(1-\tilde{\theta})^{2}\cdot\tilde{q}_{0}(1-\tilde{q}_{0})/\tilde{m}_{0}+\tilde{\theta}^{2}\cdot\tilde{q}_{1}(1-\tilde{q}_{1})/\tilde{m}_{1}}{(\tilde{q}_{0}+\tilde{q}_{1}-1)^{2}}}$ (32)
$\propto\sqrt{(1-\tilde{\theta})^{2}\cdot\tilde{q}_{0}(1-\tilde{q}_{0})/\tilde{m}_{0}+\tilde{\theta}^{2}\cdot\tilde{q}_{1}(1-\tilde{q}_{1})/\tilde{m}_{1}}$ (33)
$\propto\sqrt{(\tilde{q}_{1}-\tilde{p})^{2}\cdot\tilde{q}_{0}(1-\tilde{q}_{0})/\tilde{m}_{0}+(\tilde{p}+\tilde{q}_{0}-1)^{2}\cdot\tilde{q}_{1}(1-\tilde{q}_{1})/\tilde{m}_{1}}.$ (34)
By the arithmetic–geometric mean inequality, the minimum condition is satisfied when
$\frac{\tilde{m}_{0}}{\tilde{m}_{1}}$ $=\frac{|\tilde{q}_{1}-\tilde{p}|}{|\tilde{p}+\tilde{q}_{0}-1|}\sqrt{\frac{\tilde{q}_{0}(1-\tilde{q}_{0})}{\tilde{q}_{1}(1-\tilde{q}_{1})}}\approx\frac{|1-\tilde{p}|}{|\tilde{p}|}\sqrt{\frac{1-\tilde{q}_{0}}{1-\tilde{q}_{1}}}=(1/\tilde{p}-1)\sqrt{\kappa},$ (35)
where the approximation holds under the assumption that $\tilde{q}_{0}$ and $\tilde{q}_{1}$ are close to $1$. ∎
Proposition B.3.
[…]
@@
From Eq. (18), as $n\to\infty$ the contribution of $\hat{p}$ vanishes. Then, we have
$\displaystyle\mathrm{Var}(\hat{\theta})=\frac{(1-\theta)^{2}\cdot\frac{q(1-q)}{m-M_{1}}+\theta^{2}\cdot\frac{q(1-q)}{M_{1}}}{(2q-1)^{2}}\approx\frac{(1-\theta)^{2}\cdot\frac{q(1-q)}{m(1-\theta)}+\theta^{2}\cdot\frac{q(1-q)}{m\theta}}{(2q-1)^{2}}=\frac{1}{m}\cdot\frac{q(1-q)}{(2q-1)^{2}},$ (38)
where the approximation follows from replacing $M_{1}/m$ with its limit $\theta$ as $m\to\infty$.
@@
$\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}.$ (39)
Solving this quadratic inequality gives
@@
Finally, note that the above interval is real-valued if and only if
$q\geq\frac{1}{2}+\frac{1}{2\sqrt{2}}.$ (41)
B.3 Equivalence between the Misclassification Estimator and Our Point Estimator
We show that the misclassification estimator $\left(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\right)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] in Section 7 is equivalent to our point estimator in Eq. (8). From the definition, we have
@@ B.3 Equivalence between the Misclassification Estimator and Our Point Estimator
which exactly matches our point estimator in Eq. (8).
@@ Appendix C Code
All code used for this paper, including a plug-in Python implementation of the introduced method for LLM-as-a-judge evaluation, is available in https://github.com/UW-Madison-Lee-Lab/LLM-judge-reporting. To make this appendix self-contained, we provide below the key functions that compute the bias-adjusted estimator and its confidence interval, corresponding to the method described in Section 4.
@@
Figure 8: Python code implementation of the adjustment method described in Section 4 that computes the bias-adjusted estimate and the $(1-\alpha)$ confidence interval for the true accuracy $\theta$. The inputs p, q0, and q1 are empirical estimates from the test and calibration datasets.
Appendix D Experimental Setup for Distribution-Shift Analysis
Here, we describe the experimental setup used to produce the distribution-shift results in Figure 5. The experimental setup follows the Monte Carlo simulation in Section 8, with the calibration-set accuracy varied.
We fix the LLM judge behavior to $(q_{0},q_{1})=(0.7,0.9)$ and the test-set accuracy to $\theta_{{\mathbb{P}}}=0.5$. The calibration-set accuracy is varied over $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$, inducing label shift while keeping $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z)$ unchanged. For each setting, we generate a test dataset of size $1000$ and a calibration dataset of size $200$ with $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z=1)=\theta_{{\mathbb{Q}}}$. Results are averaged over 10,000 Monte Carlo replications.
Appendix E Additional Results on Monte Carlo Simulation
To complement the main simulation results presented in Figure 6, we report an extensive set of Monte Carlo experiments conducted across multiple configurations of the test dataset size $n=1000$, the calibration sizes $m\in\{200,500\}$, and the judge reliability parameters $(q_{0},q_{1})\in\{(0.9,0.9),(0.7,0.7),(0.9,0.7),(0.7,0.9)\}$. The remaining aspects of the simulation design follow the same setup as in the main text.
Across all combinations of $(n,m,q_{0},q_{1})$, the qualitative findings observed in the main simulation persist. Bias correction consistently improves estimation accuracy, empirical coverage attains the nominal level, and optimized calibration allocation yields shorter confidence intervals.
Below we present the complete collection of results. Each figure corresponds to one configuration $(n,m,q_{0},q_{1})$ and includes three subplots.
E.1 Results for $n=1000$ and $m=200$

### v3

@@ Acknowledgements
This paper advances the field of machine learning by improving the statistical reliability of LLM-as-a-judge evaluations. By correcting bias from imperfect LLM evaluators and providing uncertainty quantification, our work helps reduce the risk of misleading performance claims and supports more rigorous evaluation practices. The contribution is methodological and does not introduce new models that affect users but instead strengthens existing evaluation pipelines. We view the broader societal impact as positive.
Appendix A Adaptive Strategy for Allocating Calibration Samples Across Labels
We propose an adaptive strategy for allocating calibration sample sizes across the two true-label types (‘correct’ and ‘incorrect’) using a small pilot calibration set. Leveraging pilot estimates of sensitivity $\tilde{q}_{1}$ and specificity $\tilde{q}_{0}$, Algorithm 1 attains an optimal allocation under a fixed total calibration budget, leading to shorter confidence intervals.
Algorithm 1 Adaptive strategy for allocating calibration sample sizes across true label types
@@ B.1 Deriving the Variance of Estimators
We now derive the asymptotic variance of $\hat{\theta}$ using the delta method (Dorfman, 1938; Ver Hoef, 2012) for $\hat{\theta}=\frac{\hat{p}+\hat{q}_{0}-1}{\hat{q}_{0}+\hat{q}_{1}-1}$. The first order derivatives with respect to $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$ are
@@ B.2 Proofs of Propositions
Suppose that $m:=2m_{0}=2m_{1}$ and that $q:=q_{0}=q_{1}$ with $0.5<q\leq 1$. For sufficiently large $m\gtrsim 2q/(2q-1)^{2}$, the absolute bias of $\hat{\theta}$ in (2) is always smaller than that of $\hat{p}$ in (3) for all $\theta\in[0,1]$.
@@
Hence, for sufficiently large $m$ satisfying $m\gtrsim 2q/(2q-1)^{2}$, we conclude the following for all $\theta\in[0,1]$:
@@
Suppose that $n\to\infty$ and $q:=q_{0}=q_{1}$ with $\frac{1}{2}+\frac{1}{2\sqrt{2}}<q\leq 1$. Let $M_{1}\sim\mathrm{Binomial}(m,\theta)$ denote the number of ‘correct’ responses by human evaluation, and define $\hat{\phi}:=M_{1}/m$. Let $\hat{\theta}$ be the bias-corrected estimator in (11). Fix $\delta\in(0,1)$ and define $\epsilon:=\sqrt{\tfrac{\log(2/\delta)}{2m}}$. If $\epsilon<\min\{\theta,1-\theta\}$, then with probability at least $1-\delta$,
$\displaystyle\mathrm{Var}(\hat{\theta})\leq\displaystyle\mathrm{Var}(\hat{\phi})\quad\text{whenever}\quad\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}\cdot\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right).$ (30)
Moreover, since $\epsilon=O(m^{-1/2})$, as $m\to\infty$ the sufficient condition in (30) reduces to the necessary and sufficient condition $\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}$, which is equivalent to
@@
From (15), as $n\to\infty$ the contribution of $\hat{p}$ vanishes, and hence
$\displaystyle\mathrm{Var}(\hat{\theta})=\frac{(1-\theta)^{2}\cdot\frac{q(1-q)}{m-M_{1}}+\theta^{2}\cdot\frac{q(1-q)}{M_{1}}}{(2q-1)^{2}}.$ (33)
Since $\hat{\phi}=M_{1}/m$, Hoeffding’s inequality implies that for any $\delta\in(0,1)$,
$\mathop{\rm Pr}\nolimits\Big(|\hat{\phi}-\theta|\geq\epsilon\Big)\leq\delta,\qquad\epsilon:=\sqrt{\frac{\log(2/\delta)}{2m}}.$ (34)
Hence, with probability at least $1-\delta$, we have $|\hat{\phi}-\theta|\leq\epsilon$. This implies $M_{1}\geq m(\theta-\epsilon)$ and $m-M_{1}\geq m(1-\theta-\epsilon)$, and therefore
$\frac{1}{M_{1}}\leq\frac{1}{m(\theta-\epsilon)},\qquad\frac{1}{m-M_{1}}\leq\frac{1}{m(1-\theta-\epsilon)}.$ (35)
Substituting these bounds into (33) gives
@@
$=\frac{1}{m}\cdot\frac{q(1-q)}{(2q-1)^{2}}\left(\frac{(1-\theta)^{2}}{1-\theta-\epsilon}+\frac{\theta^{2}}{\theta-\epsilon}\right)$ (37)
$=\frac{1}{m}\cdot\frac{q(1-q)}{(2q-1)^{2}}\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right).$ (38)
[…]
$\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}\cdot\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right),$ (39)
which proves (16). Finally, as $m\to\infty$, we have $\epsilon=\sqrt{\log(2/\delta)/(2m)}\to 0$, and the above condition approaches $\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}$, which is then necessary and sufficient. Solving this quadratic inequality gives
@@
The interval is real-valued if and only if $q\geq\frac{1}{2}+\frac{1}{2\sqrt{2}}$. ∎
B.3 Equivalence between the Misclassification-Adjusted Estimator and Our Estimator
We show that the misclassification-adjusted estimator $\left(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\right)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] in Section 8 is equivalent to our estimator $\hat{\theta}$ in (7). From the definition, we have
@@ B.3 Equivalence between the Misclassification-Adjusted Estimator and Our Estimator
which exactly matches $\hat{\theta}$ in (7).
@@ Appendix C Code
All code used for this paper, including a plug-in Python implementation of the introduced method for LLM-as-a-judge evaluation, is available in https://github.com/UW-Madison-Lee-Lab/LLM-judge-reporting. To make this appendix self-contained, we provide below the key functions that compute the bias-corrected estimator and its confidence interval, corresponding to the method described in Section 4.
@@
Figure 8: Python code implementation of the adjustment method described in Section 4 that computes the bias-corrected estimate and the $(1-\alpha)$ confidence interval for the true accuracy $\theta$. The inputs p, q0, and q1 are empirical estimates from the test and calibration datasets.
Appendix D Additional Results on Monte Carlo Simulation
To complement the main simulation results in Figure 5, we report additional Monte Carlo experiments across multiple configurations with test set size $n=1000$, calibration sizes $m\in\{200,500\}$, and $(q_{0},q_{1})\in\{(0.9,0.9),(0.7,0.7),(0.9,0.7),(0.7,0.9)\}$. All other aspects of the simulation design follow the main-text setup.
Across all $(n,m,q_{0},q_{1})$ configurations, the main simulation findings remain consistent: bias correction improves estimation accuracy, empirical coverage matches the nominal level, and optimized calibration allocation produces shorter confidence intervals. Each figure corresponds to a specific $(n,m,q_{0},q_{1})$ setting and contains three subplots.
D.1 Results for $n=1000$ and $m=200$
@@ D.2 Results for $n=1000$ and $m=500$
Appendix E Experimental Setup for Distribution-Shift Analysis
We describe the experimental setup used to produce the distribution-shift results in Figure 7. The setup matches the Monte Carlo simulation in Section 7, except that we vary the true accuracy of the calibration dataset.
We fix the LLM judge behavior to $(q_{0},q_{1})=(0.7,0.9)$ and the test-set accuracy to $\theta_{{\mathbb{P}}}=0.5$, while varying the calibration-set accuracy over $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$. For each setting $(\theta_{{\mathbb{P}}},\theta_{{\mathbb{Q}}})$, we generate a test dataset of size $1000$ with $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z=1)=\theta_{{\mathbb{P}}}$ and a calibration dataset of size $200$ with $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z=1)=\theta_{{\mathbb{Q}}}$. We average results over 10,000 Monte Carlo replications.